OpenStudy (anonymous):

Choose the correct description of the graph of the inequality x - 4 >= 5x. A. Open circle on -1, shading to the left. B. Closed circle on -1, shading to the left. C. Open circle on -1, shading to the right. D. Closed circle on -1, shading to the right Is it D?

OpenStudy (anonymous):

Choose the correct description of the graph of the inequality -3x > 15. A. Open circle on -5, shading to the left. B. Closed circle on 5, shading to the left. C. Open circle on 5, shading to the right. D. Closed circle on -5, shading to the right. And would this be A?

OpenStudy (anonymous):

\[x-4-5x \ge 0\] \[-4x \ge 4\] then: \[x \le {4 \over -4}\] \[x \le -1\] B option

OpenStudy (anonymous):

A. Open circle on -5, shading to the left.

OpenStudy (mertsj):

The first one is B Second one is A

OpenStudy (anonymous):

What about this: Solve the inequality 6x - 8 > 4x + 10. A. x > -9 B. x > -17 C. x > 9 D. x > 17 This one i'm stuck on...

OpenStudy (anonymous):

\[6x-4x > 10+8\] \[2x > 18\] C. x > 9